# Improved Noise Schedule for Diffusion Training

### v1

@@ 3.2 Comparison with baselines and loss weight designs
We also provide a comparison with methods that adjust the loss weight, including Min-SNR and Soft-Min-SNR. We find that although these methods can achieve better results than the baseline, they are still not as effective as our method of modifying the noise schedule. This indicates that deciding where to allocate more computational resources is more efficient than adjusting the loss weight. Compared with other noise schedules like EDM Karras et al. (2022) and Flow Lipman et al. (2022), we found that no matter which CFG value, our results significantly surpass theirs under the same training iterations.
@@ 3.3 Robustness on different predicting targets
We evaluate the effectiveness of our designed noise schedule across three commonly adopted prediction targets: $\bm{\epsilon}$, $\mathbf{x}_{0}$ and $\mathbf{v}$. The results are shown in Table 4.
@@
Table 4: Effectiveness evaluated using FID-10K score on different predicting targets. The proposed Laplace schedule performs better than the baseline Cosine schedule along with training iterations.
@@ 3.5 Ablation Study
Laplace distribution is easy to implement and we adjust the scale to make the peak at the middle timestep. We conduct experiments with different Laplace distribution scales $b\in\{0.25,0.5,1.0,2.0,3.0\}$. The results are shown in Figure 3. The baseline with standard cosine schedule achieves FID score of 17.79 with CFG=1.5, 10.85 with CFG=2.0, and 11.06 with CFG=3.0 after 500K iterations. We can see that the model with Laplace distribution scale $b=0.5$ achieves the best performance 7.96 with CFG=3.0, which is relatively 26.6% better than the baseline.
Figure 3: FID-10K results on ImageNet-256 with different Laplace distribution scales $b$ in $\{0.25,0.5,1.0,2.0,3.0\}$. The location parameter $\mu$ is fixed to 0. Baseline denotes standard cosine schedule.
@@
Cosine Shifted Hoogeboom et al. (2023) is the shifted version of the standard cosine schedule. We evaluate the schedules with both positive and negative $\mu$ values. Shifted with $\mu=1$ achieves FID-10k score $\{19.34,11.67,11.13\}$ with CFG $\{1.5,2.0,3.0\}$. Results with shifted value $\mu=-1$ are $\{19.30,11.48,11.28\}$. Comparatively, both scenarios demonstrate inferior performance relative to the baseline cosine schedule ($\mu=0$). Additionally, by examining the data presented in Table 6, we find concentrated on $\lambda=0$ can best improve the results.
Cosine Scaled is also a modification of Cosine schedule. When $s$ is equal to 1, it becomes the standard Cosine version. $s>1$ means sampling more heavily around $\lambda=0$ while $s<1$ means sampling more uniformly of all $\lambda$. We report related results in Table 7. Larger values of $s(s>1)$ outperform the baseline; however, $s$ should not be excessively large and must remain within a valid range. A model trained with $s=2$ attains a score of 8.04, representing a $\mathbf{25.9\%}$ improvement over the baseline.
@@ 4 Related Work
Generally speaking, the diffusion model uses a network with shared parameters to denoise different noise intensities. However, the different noise levels may introduce conflicts during training, which makes the convergence slow. Min-SNR Hang et al. (2023) seeks the Pareto optimal direction for different tasks, achieves better convergence on different predicting targets. HDiT Crowson et al. (2024) propose a soft version of Min-SNR to further improve the efficiency on high resolution image synthesis. Stable Diffusion 3 Esser et al. (2024) puts more weight on the middle timesteps by multiplying the distribution of logit normal distribution. On the other hand, architecture modification is also explored to improve diffusion training. DiT Peebles & Xie (2023) proposes adaptive Layer Normalization with zero initialization to improve the training of Transformer architectures. A more robust ADM UNet with better training dynamics is proposed in EDM2 Karras et al. (2024) by preserving activation, weight, and update magnitudes.
@@
The design of the noise schedule plays a critical role in training diffusion models. In DDPM, Ho et al. (2020) propose linear schedule for the noise level, which is later used in Stable Diffusion Rombach et al. (2022) version 1.5 and 2.0. iDDPM Nichol & Dhariwal (2021) introduces a cosine schedule aimed at bringing the sample with the highest noise level closer to pure Gaussian noise. EDM Karras et al. (2022) proposes a new continuous framework and make the logarithm of noise intensity sampled from a Gaussian distribution. Flow matching with optimal transport Lipman et al. (2022); Liu et al. (2022) linearly interpolates the noise and data point as the input of flow-based models.  Chen (2023) underscored the need for adapting the noise schedule according to the token length, and several other works Lin et al. (2024); Tang et al. (2023) emphasize that it’s important to prevent signal leakage in the final step.
@@ 5 Conclusion
In this technical report, we present a novel method for enhancing diffusion model training by redefining the noise schedule. We theoretically analyzed that this approach equates to performing importance sampling on the noise. Empirical results show that our proposed Laplace noise schedule, focusing computational resources on mid-range steps, yields superior performance compared to the adjustment of loss weights under constrained budgets. This study not only contributes significantly to developing efficient training techniques for diffusion models but also offers potential for future large-scale applications.
Appendix A: Detailed Implementation for Noise Schedule
@@
Appendix B: Details for Sampling Process
As we mentioned before, choosing which noise schedule for sampling worth exploration. In this paper, we focus on exploring what kind of noise schedule is needed for training. Therefore, we adopted the same inference strategy as the cosine schedule to ensure a fair comparison. Specifically, first we sample $\{t_{0},t_{1},\ldots,t_{s}\}$ from uniform distribution $\mathcal{U}[0,1]$, then get the corresponding SNRs from Cosine schedule: $\{\frac{\alpha^{2}_{t_{0}}}{\sigma^{2}_{t_{0}}},\frac{\alpha^{2}_{t_{1}}}{\sigma^{2}_{t_{1}}},\ldots,\frac{\alpha^{2}_{t_{s}}}{\sigma^{2}_{t_{s}}}\}$. According to Equation 6, we get the corresponding $\{t^{\prime}_{0},t^{\prime}_{1},\ldots,t^{\prime}_{s}\}$ by inverting these SNR values through the respective noise schedules. Finally, we use DDIM Song et al. (2021) to sample with these new calculated $\{t^{\prime}\}$.

### v2

@@ 3.2 Comparison with baseline schedules and loss weight designs
We also provide a comparison with methods that adjust the loss weight, including Min-SNR and Soft-Min-SNR. Unless otherwise specified, the hyperparameter $\gamma$ for both loss weighting schemes is set to 5. We find that although these methods can achieve better results than the baseline, they are still not as effective as our method of modifying the noise schedule. This indicates that deciding where to allocate more computational resources is more efficient than adjusting the loss weight. Compared with other noise schedules like EDM Karras et al. (2022) and Flow Matching Lipman et al. (2022), we found that no matter which CFG value, our results significantly surpass theirs under the same training iterations.
@@ 3.3 Robustness on different predicting targets
We evaluate the effectiveness of our designed noise schedule across three commonly adopted prediction targets: $\bm{\epsilon}$, $\mathbf{x}_{0}$, and $\mathbf{v}$. The results are shown in Table 4.
@@
Table 4: Effectiveness evaluated using FID-10K score on different predicting targets: $\mathbf{x}_{0}$, $\epsilon$, and $\mathbf{v}$. The proposed Laplace schedule performs better than the baseline Cosine schedule along with training iterations.
@@ 3.5 Ablation Study
Laplace distribution, known for its simplicity and exponential decay from the center, is straightforward to implement. We leverage its symmetric nature and adjust the scale parameter to center the peak at the middle timestep. We conduct experiments with different Laplace distribution scales $b\in\{0.25,0.5,1.0,2.0,3.0\}$. The results are shown in Figure 3. The baseline with standard cosine schedule achieves FID score of 17.79 with CFG=1.5, 10.85 with CFG=2.0, and 11.06 with CFG=3.0 after 500K iterations. We can see that the model with Laplace distribution scale $b=0.5$ achieves the best performance 7.96 with CFG=3.0, which is relatively 26.6% better than the baseline.
Figure 3: FID-10K results on ImageNet-256 with location parameter $\mu$ fixed to 0 and different Laplace distribution scales $b$ in $\{0.25,0.5,1.0,2.0,3.0\}$. Baseline denotes standard cosine schedule.
@@
Cosine Shifted Hoogeboom et al. (2023) is the shifted version of the standard cosine schedule. We evaluate the schedules with both positive and negative $\mu$ values to comprehensively assess its impact on model performance. Shifted with $\mu=1$ achieves FID-10k score $\{19.34,11.67,11.13\}$ with CFG $\{1.5,2.0,3.0\}$. Results with shifted value $\mu=-1$ are $\{19.30,11.48,11.28\}$. Comparatively, both scenarios demonstrate inferior performance relative to the baseline cosine schedule ($\mu=0$). Additionally, by examining the data presented in Table 6, we find concentrated on $\lambda=0$ can best improve the results.
Cosine Scaled is also a modification of Cosine schedule. When $s=1$, it becomes the standard Cosine version. $s>1$ means sampling more heavily around $\lambda=0$ while $s<1$ means sampling more uniformly of all $\lambda$. We report related results in Table 7. Our experimental results reveal a clear trend: larger values of $s(s>1)$ consistently outperform the baseline, highlighting the benefits of focused sampling near $\lambda=0$. However, it’s crucial to note that $s$ should not be excessively large and must remain within a valid range to maintain stable training dynamics. For example, decreasing $1/s$ from 0.5 to 0.25 hurts the performance and cause the FID score to drop. Striking the right balance is key to optimizing performance. In our experiments, a model trained with $s=2$ achieved a remarkable score of 8.04, representing a substantial $\mathbf{25.9\%}$ improvement over the baseline.
The experiments with various noise schedules, including Laplace, Cauchy, Cosine Shifted, and Cosine Scaled, reveal a shared phenomenon: models perform better when the noise distribution or schedule is concentrated around $\lambda=0$. For the Laplace distribution, a scale of $b=0.5$ yielded the best performance, outperforming the baseline by 26.6%. In the case of the Cauchy distribution, models with a location parameter $\mu=0$ performed better than those with $\mu$ values biased towards negative or positive directions. The Cosine Shifted schedule showed inferior performance when shifted away from $\mu=0$, while the Cosine Scaled schedule demonstrated that larger values of $s$ (sampling more heavily around $\lambda=0$) consistently outperformed the baseline, with an optimal improvement of 25.9% at $s=2$. This consistent trend suggests that focusing the noise distribution or schedule near $\lambda=0$ is beneficial for model performance. While these different schedules take various mathematical forms, they all achieve similar optimal performance when given equivalent training budgets. The specific mathematical formulation is less crucial than the underlying design philosophy: increasing the sampling probability of intermediate noise levels. This principle provides a simple yet effective guideline for designing noise schedules.
@@ 4 Related Works
Generally speaking, the diffusion model uses a network with shared parameters to denoise different noise intensities. However, the different noise levels may introduce conflicts during training, which makes the convergence slow. P2 Choi et al. (2022) improves image generation performance by prioritizing the learning of perceptually rich visual concepts during training through a redesigned weighting scheme. Min-SNR Hang et al. (2023) seeks the Pareto optimal direction for different tasks, achieves better convergence on different predicting targets. HDiT Crowson et al. (2024) propose a soft version of Min-SNR to further improve the efficiency on high resolution image synthesis. Stable Diffusion 3 Esser et al. (2024) puts more sampling weight on the middle timesteps by multiplying the distribution of logit normal distribution.
On the other hand, architecture modification is also explored to improve diffusion training. DiT Peebles & Xie (2023) proposes adaptive Layer Normalization with zero initialization to improve the training of Transformer architectures. Building upon this design, MM-DiT Esser et al. (2024) extends the approach to a multi-modal framework (text to image) by incorporating separate sets of weights for each modality. HDiT Crowson et al. (2024) uses a hierarchical transformer structure for efficient, linear-scaling, high-resolution image generation. A more robust ADM UNet with better training dynamics is proposed in EDM2 Karras et al. (2024) by preserving activation, weight, and update magnitudes. In this work, we directly adopt the design from DiT Peebles & Xie (2023) and focus on investigating the importance sampling schedule in diffusion models.
@@
The design of the noise schedule plays a critical role in training diffusion models. In DDPM, Ho et al. (2020) propose linear schedule for the noise level, which was later adopted by Stable Diffusion Rombach et al. (2022) version 1.5 and 2.0. However, the linear noise schedule introduces signal leakage at the highest noise step Lin et al. (2024); Tang et al. (2023), hindering performance when sampling starts from a Gaussian distribution. Improved DDPM Nichol & Dhariwal (2021) introduces a cosine schedule aimed at bringing the sample with the highest noise level closer to pure Gaussian noise. EDM Karras et al. (2022) proposes a new continuous framework and make the logarithm of noise intensity sampled from a Gaussian distribution. Flow matching with optimal transport Lipman et al. (2022); Liu et al. (2022) linearly interpolates the noise and data point as the input of flow-based models.  Chen (2023) underscored the need for adapting the noise schedule according to the image resolution. Hoogeboom et al. (2023) found that cosine schedule exhibits superior performance for images of $32\times 32$ and $64\times 64$ resolutions and propose to shift the cosine schedule to train on images with higher resolutions.
@@ 5 Conclusion
In this paper, we present a novel method for enhancing the training of diffusion models by strategically redefining the noise schedule. Our theoretical analysis demonstrates that this approach is equivalent to performing importance sampling on the noise. Empirical results show that our proposed Laplace noise schedule, which focuses computational resources on mid-range noise levels, yields superior performance compared to adjusting loss weights under constrained computational budgets. This study not only contributes significantly to the development of efficient training techniques for diffusion models but also offers promising potential for future large-scale applications.
Appendix A Appendix
A.1 Detailed Implementation for Noise Schedule
@@
A.2 Details for proposed Laplace and Cauchy Design
For a Laplace distribution with location parameter $\mu$ and scale parameter $b$, the probability density function (PDF) is given by:
$p(\lambda)=\frac{1}{2b}\exp\left(-\frac{|\lambda-\mu|}{b}\right)$ (7)
The cumulative distribution function (CDF) can be derived as follows:
$1-t$ $=\int_{-\infty}^{\lambda}p(x)\,\mathrm{d}x$
$=\int_{-\infty}^{\lambda}\frac{1}{2b}\exp\left(-\frac{|x-\mu|}{b}\right)\,\mathrm{d}x$
$=\frac{1}{2}\left(1+\text{sgn}(\lambda-\mu)\left(1-\exp\left(-\frac{|\lambda-\mu|}{b}\right)\right)\right)$
To obtain $\lambda$ as a function of $t$, we solve the inverse function:
$\lambda$ $=\mu-b\text{sgn}(0.5-t)\ln(1-2|t-0.5|)$
For a Cauchy distribution with location parameter $\mu$ and scale parameter $\gamma$, the PDF is given by:
$f(\lambda;\mu,\gamma)=\frac{1}{\pi\gamma}\left[1+\left(\frac{\lambda-\mu}{\gamma}\right)^{2}\right]^{-1}$ (8)
The corresponding CDF is:
$F(\lambda;\mu,\gamma)=\frac{1}{2}+\frac{1}{\pi}\arctan\left(\frac{\lambda-\mu}{\gamma}\right)$ (9)
To derive $\lambda(t)$, we proceed as follows:
$1-t$ $=F(\lambda;\mu,\gamma)$ (10)
$1-t$ $=\frac{1}{2}+\frac{1}{\pi}\arctan\left(\frac{\lambda-\mu}{\gamma}\right)$ (11)
$t$ $=\frac{1}{2}-\frac{1}{\pi}\arctan\left(\frac{\lambda-\mu}{\gamma}\right)$ (12)
Solving for $\lambda$, we obtain:
$\lambda(t)=\mu+\gamma\tan\left(\frac{\pi}{2}(1-2t)\right)$ (13)
A.3 Combination between noise schedule and timestep importance sampling
We observe that incorporating importance sampling of timesteps into the cosine schedule bears similarities to the Laplace schedule. Typically, the distribution of timestep $t$ is uniform $\mathcal{U}[0,1]$. To increase the sampling frequency of middle-level timesteps, we propose modifying the sampling distribution to a simple polynomial function:
$p(t^{\prime})=\begin{cases}C\cdot t^{\prime n},&t^{\prime}<\frac{1}{2}\\ C\cdot\left(1-t^{\prime}\right)^{n},&t^{\prime}\geq\frac{1}{2},\end{cases}$ (14)
where $C=(n+1)2^{n}$ is the normalization factor ensuring that the cumulative distribution function (CDF) equals 1 at $t=1$.
To sample from this distribution, we first sample $t$ uniformly from $(0,1)$ and then map it using the following function:
$t^{\prime}=\begin{cases}\left(\frac{1}{2}\right)^{\frac{n}{n+1}}t^{\frac{1}{n+1}},&t<\frac{1}{2}\\ 1-\left(\frac{1}{2}\right)^{\frac{n}{n+1}}(1-t)^{\frac{1}{n+1}},&t\geq\frac{1}{2},\end{cases}$ (15)
We incorporate the polynomial sampling of $t$ into the cosine schedule $\lambda=-2\log\tan\frac{\pi t}{2}$, whose inverse function is $t=\frac{2}{\pi}\arctan\exp\left(-\frac{\lambda}{2}\right)$. Let us first consider the situation where $t<\frac{1}{2}$:
$\left(\frac{1}{2}\right)^{\frac{n}{n+1}}t^{\frac{1}{n+1}}$ $=\frac{2}{\pi}\arctan\exp\left(-\frac{\lambda}{2}\right)$ (16)
$t$ $=2^{n}\left(\frac{2}{\pi}\arctan\exp\left(-\frac{\lambda}{2}\right)\right)^{n+1}$ (17)
We then derive the expression with respect to $\mathrm{d}\lambda$:
$\frac{\mathrm{d}t}{\mathrm{d}\lambda}$ $=2^{n}\left(\frac{2}{\pi}\right)^{n+1}(n+1)\left(\arctan\exp\left(-\frac{\lambda}{2}\right)\right)^{n}\frac{1}{1+\exp(-\lambda)}\frac{1}{-2}\exp(-\lambda/2)$ (18)
$p(\lambda)$ $=(n+1)\frac{4^{n}}{\pi^{(n+1)}}\arctan^{n}\exp\left(-\frac{\lambda}{2}\right)\frac{\exp(-\frac{1}{2}\lambda)}{1+\exp(-\lambda)}$ (19)
Considering symmetry, we obtain the final distribution with respect to $\lambda$ as follows:
$p(\lambda)$ $=(n+1)\frac{4^{n}}{\pi^{(n+1)}}\arctan^{n}\exp\left(-\frac{|\lambda|}{2}\right)\frac{\exp(-\frac{1}{2}|\lambda|)}{1+\exp(-|\lambda|)}$ (21)
We visualize the schedule discussed above and compare it with Laplace schedule in Figure 4. We can see that $b=1$ for Laplace and $n=2$ for cosine-ply matches well. We also conduct experiments on such schedule and present results in Table 8. They perform similar and both better than the standard cosine schedule.
We visualize the schedules discussed above and compare them with the Laplace schedule in Figure 4. The results demonstrate that Laplace with $b=1$ and cosine-ply with $n=2$ exhibit a close correspondence. To evaluate the performance of these schedules, we conducted experiments and present the results in Table 8. Both the Laplace and cosine-ply schedules show similar performance, and both outperform the standard cosine schedule.
Iterations 100,000 200,000 300,000 400,000 500,000
Cosine-ply ($n=2$) 28.65 13.77 10.06 8.69 7.98
Laplace ($b=1$) 28.89 13.90 10.17 8.85 8.19
Table 8: Performance comparison of cosine-ply ($n=2$) and Laplace ($\mu=1$) schedules over different iteration counts
Figure 4: Visualization of $p(\lambda)$ for Laplace schedule and cosine schedule with polynomial timestep sampling.
A.4 Flow Matching with Logit-Normal Sampling
In Stable Diffusion 3 Esser et al. (2024) and Movie Gen Polyak et al. (2024), logit-normal sampling is applied to improve the training efficiency of flow models. To better understand this approach, we present a detailed derivation from the logit-normal distribution to the probability density function of logSNR $\lambda$.
Let the Logit transformation $X=\text{logit}(t)$ of random variable $t$ follow a normal distribution:
$X\sim\mathcal{N}(\mu,\sigma^{2})$ (22)
Then, the probability density function of $t$ is:
$p(t;\mu,\sigma)=\frac{1}{\sigma\cdot t\cdot(1-t)\cdot\sqrt{2\pi}}\exp\left(-\frac{(\text{logit}(t)-\mu)^{2}}{2\sigma^{2}}\right),\quad t\in(0,1)$ (23)
where $\text{logit}(t)=\log\left(\frac{t}{1-t}\right)$, and $\mu$ and $\sigma$ are constants.
Consider the variable transformation:
$\lambda=2\log\left(\frac{1-t}{t}\right)$ (24)
Our goal is to find the probability density function $p(\lambda)$ of random variable $\lambda$.
First, we solve for $t$ in terms of $\lambda$:
$\frac{\lambda}{2}$ $=\log\left(\frac{1-t}{t}\right)$
$e^{\frac{\lambda}{2}}$ $=\frac{1-t}{t}$
$1-t$ $=te^{\frac{\lambda}{2}}$
11 11 $=t\left(1+e^{\frac{\lambda}{2}}\right)$
$t(\lambda)$ $=\frac{1}{1+e^{\frac{\lambda}{2}}}$
Next, we calculate the Jacobian determinant $\left|\frac{\mathrm{d}t}{\mathrm{d}\lambda}\right|$:
$t(\lambda)$ $=\frac{1}{1+e^{\frac{\lambda}{2}}}$
$\frac{\mathrm{d}t}{\mathrm{d}\lambda}$ $=-\frac{e^{\frac{\lambda}{2}}\cdot\frac{1}{2}}{(1+e^{\frac{\lambda}{2}})^{2}}$
$\left|\frac{\mathrm{d}t}{\mathrm{d}\lambda}\right|$ $=\frac{e^{\frac{\lambda}{2}}}{2(1+e^{\frac{\lambda}{2}})^{2}}$
Using the variable transformation formula:
$p(\lambda)=p(t(\lambda);\mu,\sigma)\cdot\left|\frac{\mathrm{d}t}{\mathrm{d}\lambda}\right|$ (25)
We calculate $p(t(\lambda);\mu,\sigma)$:
$\text{logit}(t(\lambda))$ $=\log\left(\frac{t(\lambda)}{1-t(\lambda)}\right)=\log\left(\frac{\frac{1}{1+e^{\frac{\lambda}{2}}}}{\frac{e^{\frac{\lambda}{2}}}{1+e^{\frac{\lambda}{2}}}}\right)=-\frac{\lambda}{2}$
$p(t(\lambda);\mu,\sigma)$ $=\frac{(1+e^{\frac{\lambda}{2}})^{2}}{\sigma e^{\frac{\lambda}{2}}\sqrt{2\pi}}\exp\left(-\frac{(\mu+\frac{\lambda}{2})^{2}}{2\sigma^{2}}\right)$
Multiplying by the Jacobian determinant:
$p(\lambda)$ $=\frac{(1+e^{\frac{\lambda}{2}})^{2}}{\sigma e^{\frac{\lambda}{2}}\sqrt{2\pi}}\exp\left(-\frac{(\mu+\frac{\lambda}{2})^{2}}{2\sigma^{2}}\right)\cdot\frac{e^{\frac{\lambda}{2}}}{2(1+e^{\frac{\lambda}{2}})^{2}}$
$=\frac{1}{2\sigma\sqrt{2\pi}}\exp\left(-\frac{(\lambda+2\mu)^{2}}{8\sigma^{2}}\right)$
Therefore, the probability density function of $\lambda$ is:
$p(\lambda)=\frac{1}{2\sigma\sqrt{2\pi}}\exp\left(-\frac{(\lambda+2\mu)^{2}}{8\sigma^{2}}\right),\quad\lambda\in(-\infty,+\infty)$ (26)
This shows that $\lambda$ follows a normal distribution with mean $-2\mu$ and variance $4\sigma^{2}$:
$\lambda\sim\mathcal{N}(-2\mu,4\sigma^{2})$ (27)
The mean and variance are:
$\mathbb{E}[\lambda]$ $=-2\mu$
$\text{Var}(\lambda)$ $=4\sigma^{2}$
To verify normalization, we integrate $p(\lambda)$ over its domain:
$\int_{-\infty}^{+\infty}p(\lambda)\,\mathrm{d}\lambda$ $=\int_{-\infty}^{+\infty}\frac{1}{2\sigma\sqrt{2\pi}}\exp\left(-\frac{(\lambda+2\mu)^{2}}{8\sigma^{2}}\right)\mathrm{d}\lambda$
Let $\text{Let }z$ $=\frac{\lambda+2\mu}{2\sqrt{2}\sigma}\Rightarrow\mathrm{d}\lambda=2\sqrt{2}\sigma\,\mathrm{d}z$
$=\frac{2\sqrt{2}\sigma}{2\sigma\sqrt{2\pi}}\int_{-\infty}^{+\infty}e^{-z^{2}}\mathrm{d}z$
$=\frac{1}{\sqrt{\pi}}\cdot\sqrt{\pi}=1$
Thus, $p(\lambda)$ satisfies the normalization condition for probability density functions.
We compare the standard cosine scheudle Nichol & Dhariwal (2021), Flow Matching Liu et al. (2022); Lipman et al. (2022), and Flow Matching with Logit-normal sampling Esser et al. (2024); Polyak et al. (2024). The probability density functions of these schedules are visualized in Figure 5. Our analysis reveals that Flow Matching with Logit-normal sampling concentrates more probability mass around $\lambda=0$ compared to both the standard Cosine and Flow Matching schedules, resulting in improved training efficiency Esser et al. (2024); Polyak et al. (2024).
Figure 5: Comparison of probability density functions for different flow matching approaches. The plot shows three distributions: Flow Matching with Logit-Normal sampling (blue), Flow Matching without Logit-Normal sampling (green), and the Cosine schedule (orange).
A.5 Importance of Time Intervals
To investigate the significance of training intervals, we conducted controlled experiments using a simplified setup. We divided the time range $(0,1)$ into four equal segments: $\text{bin}_{i}=\left(\frac{i}{4},\frac{i+1}{4}\right),i=0,1,2,3$. We first trained a base model $\mathbf{M}$ over the complete range $(0,1)$ for 1M iterations, then fine-tuned it separately on each bin for 140k iterations to obtain four specialized checkpoints $\mathbf{m}_{i},i=0,1,2,3$.
For evaluation, we designed experiments using both the base model $\mathbf{M}$ and fine-tuned checkpoints $\mathbf{m}_{i}$. To assess the importance of each temporal segment, we selectively employed the corresponding fine-tuned checkpoint during its specific interval while maintaining the base model for remaining intervals. For example, when evaluating $\text{bin}_{0}$, we used $\mathbf{m}_{0}$ within its designated interval and $\mathbf{M}$ elsewhere.
The FID results across these four experimental configurations are presented in Figure 6. Our analysis reveals that optimizing intermediate timesteps (bin1 and bin2) yields superior performance, suggesting the critical importance of these temporal regions in the diffusion process.
Figure 6: Comparative analysis of interval-specific fine-tuning effects. When sampling within interval $\left(\frac{1}{4},\frac{2}{4}\right)$, “Bin1” indicates the use of fine-tuned weights $\mathbf{m}_{1}$, while $\mathbf{M}$ is used for other intervals. “Baseline” represents the use of base model $\mathbf{M}$ throughout all intervals, and “All Tuned” denotes the application of interval-specific fine-tuned models within their respective ranges.
A.6 Importance Sampling as Loss Weight
We investigate the comparative effectiveness of our approach when applied as a noise schedule versus a loss weighting mechanism. We adopt Equation 21 as our primary noise schedule due to its foundation in the cosine schedule and demonstrated superior FID performance. To evaluate its versatility, we reformulate the importance sampling as a loss weighting strategy and compare it against established weighting schemes, including Min-SNR and Soft-Min-SNR.
Cosine Cosine-Ply ($n$=2) Min-SNR Soft-Min-SNR Cosine-Ply as weight
FID-10K 10.85 7.98 9.70 9.07 8.88
Table 9: Quantitative comparison of different noise scheduling strategies and loss weighting schemes. Lower FID scores indicate better performance.
Figure 7 illustrates the loss weight derived from Cosine-Ply ($n$=2) schedule alongside Min-SNR and Soft-Min-SNR. We can observe that under the setting of predict target as $\mathbf{v}$, Min-SNR and Soft-Min-SNR can be seemed as putting more weight on intermediate levels, aligning with our earlier findings on the importance of middle-level noise densities.
Figure 7: Visualization of different loss weight schemes.
A.7 Additional experiments on other datasets
ImageNet, comprising over one million natural images, has been widely adopted as a benchmark dataset for validating improvements in diffusion models Peebles & Xie (2023); Karras et al. (2024).
In addition to ImageNet, we evaluate our approach on the CelebA Liu et al. (2015) dataset ($64\times 64$ resolution in pixel space), which consists of face images. We employ a DiT architecture (12 layers, embedding dimension of 512, 8 attention heads, and patch size of 4) using different noise schedules. This is an unconditional generation setting within a single domain. We present FID results as follows:
FID $\downarrow$ 100k 150k
cosine 10.0696 7.93795
Laplace (ours) 7.93795 6.58359
Table 10: FID scores on CelebA dataset at different training iterations
We also follow Stable Diffusion 3 Esser et al. (2024), train on a more complicated dataset CC12M Changpinyo et al. (2021) dataset (over 12M image-text pairs) and report the FID results here. We download the dataset using webdataset. We train a DiT-base model using CLIP as text conditioner. The images are cropped and resized to $256\times 256$ resolution, compressed to $32\times 32\times 4$ latents and trained for 200k iterations at batch size 256.
FID $\downarrow$ 200k
cosine 58.3619
Laplace (ours) 54.3492 (-4.0127)
Table 11: FID scores on CC12M dataset at 200k iterations
Our method demonstrated strong generalization capabilities across both unconditional image generation using the CelebA dataset and text-to-image generation using the CC12M dataset.
A.8 Additional Visual Results
We present addition visual results in Figure 8 to demonstrate the differences in generation quality between models trained with Cosine and our proposed Laplace schedule. Each case presents two rows of outputs, where the upper row shows results from the cosine schedule and the lower row displays results from our Laplace schedule. Each row contains five images corresponding to models trained for 100k, 200k, 300k, 400k, and 500k iterations, illustrating the progression of generation quality across different training stages. For each case, the initial noise inputs are identical. As shown in the results, our method achieves faster convergence in both basic object formation (at 100k iterations) and fine detail refinement, demonstrating superior learning efficiency throughout the training process.
Figure 8: Visual comparison of results generated by model trained by cosine schedule and our proposed Laplace. For each case, the above row is generated by cosine schedule, the below is generated by Laplace. The 5 images from left to right represents the results generated by the model trained for 100k, 200k, 300k, 400k, and 500k iterations.